# Neural Networks for Tamed Milstein Approximation of SDEs with Additive Symmetric Jump Noise Driven by a Poisson Random Measure

### v1

@@ I Introduction
Stochastic differential equations (SDEs) are essential tools for modeling systems driven by uncertainty and noise. In finance, the well-known Vasicek and Black-Scholes models describe, respectively, the evolution of interest rates and the pricing of European options via diffusion-type dynamics Vasicek (1977); Black and Scholes (1973). However, many phenomena exhibit abrupt changes that standard continuous models fail to capture. To address this, jump-diffusion models such as those introduced by Merton (1976), Kou (2002), and Duffie et al. (2000) incorporate sudden discontinuities into the stochastic framework, allowing for more realistic modeling of asset prices, risk, and volatility. Beyond finance, SDEs with or without jumps have been used in diverse fields including ecology, epidemiology, neuroscience, and climate science, where they effectively model population dynamics Marrec et al. (2023), disease propagation Liu et al. (2021), neuronal variability Faisal et al. (2008), and climate fluctuations Hasselmann (1976). These broad applications motivate the study and development of general stochastic models and inference techniques capable of handling both continuous and discontinuous sources of randomness.
@@ Appendix B Density estimation
Inference for Lévy processes has been developed through parametric, nonparametric, spectral, and Bayesian methods. Masuda Masuda (2015) presents a detailed study of parametric estimation for discretely observed Lévy processes, focusing on quasi-likelihood methods and their asymptotic properties under infinite activity. In the nonparametric setting, Figueroa-López and Houdré Figueroa-López and Houdré (2006) derive non-asymptotic risk bounds for Lévy density estimators using model selection techniques, while Figueroa-López Figueroa-López (2011) proposes sieve-based estimators with pointwise and uniform confidence bands. Belomestny Belomestny (2010) introduces a spectral method to estimate the fractional index of a Lévy process from its characteristic function. Fourier-based approaches are further developed by Belomestny and Reiss Belomestny and Reiss (2006), and Comte and Genon-Catalot Comte and Genon-Catalot (2009), who construct regularized estimators for the Lévy density via inversion of the characteristic function. Nickl and Reiss Nickl and Reiss (2012) provide a rigorous analysis of minimax optimality and posterior contraction rates under low- and high-frequency regimes. In the Bayesian setting, Jasra et al. Jasra et al. (2019) apply MCMC methods for stable Lévy-driven SDEs with high-frequency observations, and Belomestny et al. Belomestny et al. (2018) propose a nonparametric Bayesian estimator for Gamma-type subordinators. These works provide theoretical and computational tools for inference in a broad class of Lévy-driven models.

### v2

@@ I Introduction
Stochastic differential equations (SDEs) are essential tools for modeling systems driven by uncertainty and noise. In finance, the well-known Vasicek and Black–Scholes models describe, respectively, the evolution of interest rates and the pricing of European options via diffusion-type dynamics Vasicek (1977); Black and Scholes (1973). However, many phenomena exhibit abrupt changes that standard continuous models fail to capture. To address this, jump-diffusion models such as those introduced by Merton (1976); Kou (2002); Duffie et al. (2000) incorporate sudden discontinuities into the stochastic framework, allowing for more realistic modeling of asset prices, risk, and volatility. Beyond finance, SDEs with or without jumps have been used in diverse fields including ecology, neuroscience, and climate science, where they effectively model population dynamics Marrec et al. (2023), neuronal variability Faisal et al. (2008), and climate fluctuations Hasselmann (1976). These broad applications motivate the study and development of general stochastic models and inference techniques capable of handling both continuous and discontinuous sources of randomness.
@@ Appendix B Density estimation
Inference for Lévy processes has been extensively studied using parametric, nonparametric, spectral, and Bayesian methodologies. In the parametric setting, Masuda Masuda (2015) provides a comprehensive overview of estimation techniques for discretely observed Lévy processes, focusing on quasi-likelihood methods. For nonparametric inference, Comte and Genon-Catalot Comte and Genon-Catalot (2009) propose kernel-type estimators for the Lévy density using high-frequency data under a pure-jump framework. Figueroa-López and Houdré Figueroa-López and Houdré (2006) derive non-asymptotic risk bounds for Lévy density estimators via penalized model selection, while Figueroa-López Figueroa-López (2011) constructs sieve-based estimators with provable confidence bands for the Lévy density. Belomestny Belomestny (2010) introduces a spectral method to estimate the fractional index of a Lévy process from the empirical characteristic function, and Belomestny and Reiß Belomestny and Reiß (2015) develop Fourier-based estimators for the Lévy density using regularized inversion of the characteristic function. In the Bayesian setting, Jasra et al. Jasra et al. (2019) develop a quasi-likelihood-based MCMC method for stable Lévy-driven SDEs under high-frequency observation. They prove posterior consistency and asymptotic normality via a Bernstein–von Mises theorem and propose a scaling strategy to ensure algorithmic efficiency as data frequency increases. Additionally, Belomestny et al. Belomestny et al. (2019) propose a nonparametric Bayesian method to estimate the Lévy density of infinite activity subordinators using low-frequency data. Their approach combines data augmentation via Gamma process bridges with an infinite-dimensional reversible jump MCMC algorithm. The method achieves posterior consistency and performs well on both simulated and real insurance datasets. These works provide theoretical and computational tools for inference in a broad class of Lévy-driven models.